OpenStudy (anonymous):

(4, -7); 4x - 5y = 3 Write an equation of the line containing the given point & parallel to the given line. y=

OpenStudy (anonymous):

first find the slope of the given line. I'd rewrite it in the form y = mx+b to make this easy. y = 4/(5)x-3/(5), so the slope is 4/5

OpenStudy (anonymous):

next. we can write the equation: y = 4/5(x-4) - 7

OpenStudy (anonymous):

ok I tried that answer nad it did not like it. It wants it in the form of y=mx+b

OpenStudy (anonymous):

ah, so you'll need to distribute the 4/5 through the parentheses

OpenStudy (anonymous):

I put the () on it but is says it is not fully simplified

OpenStudy (anonymous):

y = 4x/5-19/5